y varies directly as x and inversely as z. y=4/3 when x=3 and z=6. find y when x=4 and z=3
Answer by stanbon(75887)

y varies directly as x and inversely as z.
-------
y = k*x/z
------
Solve for "k" using "y=4/3 when x=3 and z=6.
4/3 = k*3/6
---
8 = 3k
k = 8/3
---------
Equation:
y = (8/3)*x/z
---------------------------
Find y when x=4 and z=3
y = (8/3)*4/3
-----------
y = 32/9 = 3 5/9
=======================
Cheers,
Stan H.
=======================
-----